OpenStudy (anonymous):

Convert 3.341463415 into a reduced, mixed number.

OpenStudy (anonymous):

\[3+\frac{341463415}{1000000000}\text{= }3+\frac{68292683}{200000000}\]From Mathematica, 341463415 is divisible by 5, 1277 and 53479

OpenStudy (anonymous):

\[3 \frac{ 14 }{ 41 }\]